What does acquisition mean in the nfl gaints helmet uno deck of cards?
"When you play this card you gain the ability to make major moves and acquire a card that can improve your hand and get you closer to winning the game. Play the card and your can trade in one card from your hand for a card from any opponent by calling out loud the card you wish to acquire. If an opponent have that card they must acknowledge it and accept whatever card you offer in trade. In the event that there is more than one opponent holding the sought after card, you have the option to trade with both or only one. If no opponent have the sought after card, then you call off the color to follow and play resumes in the direction it was proceeding." - from the rules of our Bears deck, assume it is the same.
What is the phone number of the Rolling Meadows Historical Msm in Rolling Meadows Illinois?
The phone number of the Rolling Meadows Historical Msm is: 847-577-7086.
What does it mean for two events to be dependent?
An even is dependent of the other when one has to occur before the other can do so. Else, one can also occur with the other.
E.g. Christmas (or the winter holidays if you will) has to happen/so that/you go buy Christmas presents.
Hunger usually "depends" on the lack of food in the stomach.
---
In math, events are dependent when one event affects the probability of the other.
E.g. There is 1 red, 1 green and 1 blue balls in a container. You pick a ball out of 3. There is a probability of 1/3 = 33% of being a certain colour. Now there are only two balls in the container. The next one you pick has a probability of 1/2=50% of being a certain colour,
When you roll a dice what is the probability it will land on 1?
You cannot roll "a dice" because it is one die, many dice.
If you roll an ordinary, 6 faced die, the probability that it will land on 1 is 1/6.
Drawing a seven from a deck of cards?
There is a 1 in 52 chance that the first card drawn in a standard, shuffled 52 card deck, will have a value of 7.
What is a probability model with infinite sample space?
The time taken between your posting this question and receiving a reply.
What is a square number less than 16 between 1-40 with a probability as a fraction?
9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.
9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.
9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.
9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.
I would think about 70% of them need to be correct
On a single blind draw, it's 4/20 = 20% .
What does it mean low correlation to other assets?
In statistics, correlation measures how strong the relationship is between two entities. A correlation of 1.0 means that the two move in perfect tandem with each other. A correlation of zero means that the relationship between them is totally random. A negative correlation, unusual in the investing world, means that they move in opposite directions. In investing, low correlation means that different asset types have not performed in the same way: When returns on some asset types were declining, returns on others were declining less, or indeed gaining. For investors, this diversification has obvious benefits: If poor performance in one investment can be offset by better (or even good) performance in another, extreme losses in an overall portfolio will be rarer than otherwise, and the capital will grow more in the long run.
Is it still possible to play a game of uno with three number cards missing from a deck?
you can play, just make sure to notify everyone to the missing cards
When rolling a fair dice numbered one to six what is the probability of rolling a one?
The probability of rolling any single number of a regular die on one roll is one in six, or 1/6, or 0.166666....
What is the probability that the letter chosen will be a t?
The answer depends on the context. From the English alphabet, it is 1/26 = 0.038. In a typical English language document it is approx 0.091. From a set of Scrabbles tiles it is 0.06. From my first name it is 0.
in animals -- build bones. It is in many useful building materials. Concrete, limestone in many forms,
If you spun a coin 5 times how likely would it be to land on heads every time?
The probability of a spinning coin landing on heads is 50% each time.
The probability of five consecutive heads is (0.5 x 0.5 x 0.5 x 0.5 x 0.5) = 1/32 = 0.03125 = (3 and 1/8)%
What is the importance of the concept of probability?
An accurate knowledge of the probability of some event to occur allows the concepts of planning, preparing and expecting to take on meaning. If you had absolutely no idea what was going to happen from one moment to the next, it would be completely illogical for you to do any of those above things. Not to mention that the entire idea of "surprise" would be moot since everything would be a surprise.
No. Probabilities are 'counted' as between 1 (certain) and 0 (impossible)
What is the probability of getting 2 heads if a coin is tossed thrice?
(1/2)^3 = 1/8th
Since the initial probability (assuming independence) of getting a head in a single toss is one half (1/2) we just cube this probability because of the number of events we are performing. So if you were to try to calculate the probability of a coin being tossed 6 times it would be one half to the 6th power which is 1/64.
A pair of dice is rolled What is the probability that the sum is equal to 5?
Number of possible total pairs = 36
Number of pairs which give five as a sum = {1,4},{2,3},{3,2},{4,1}
So, P(Sum=5) = 4/36 = 1/9